2021 was 50,000. Compute selected ratios, and compare liquidity, profitability, and solvency for two companies P15.5 (LO 2) Selected financial data of Target (USA) and Wal-Mart Stores, Inc. (USA) for a recent year are presented below (in millions). Wal-Mart Target Stores, Inc. Corporation Income Statement Data for Year $476,294 358,069 91,353 $72,596 51,160 Net sales Cost of goods sold 16,816 Selling and administrative expenses Interest expense Other income (expense) Income tax expense 2,335 1,126 (410) 8,105 (391) 1,132 $ 16,022 $ 1,971 000 EC Net income Statement of Financial Position 00 08 Data (End of Year) $32,980 11,573 $143,566 Non-current assets 61,185 Current assets $204,751 $44,553 Total assets $ 81,339 54,067 69,345 $16,231 Total equity 15,545 Non-current liabilities 12,777 Current liabilities $204,751 $44,553 Total equity and liabilities Beginning-of-Year Balances $203,105 81,738 $48,163 Total assets 16,558 Total equity Current liabilities 71,818 14,031 31,605 131,287 Total liabilities Other Data $ 6,723 44,331 $2,921 Average net accounts receivable Average inventory Net cash provided by operating activities 8,335 23,257 6,520 Instructions a. For each company, compute the following ratios. 1. Current ratio. 2. Accounts receivable turnover.
3. Average collection period. a. 4. Inventory turnover Target 6.1 4. Inventory turnover 5. Days in inventory. 6. Profit margin. 7. Asset turnover. 8. Return on assets. 9. Return on ordinary shareholders' equity. 10. Debt to assets ratio. 11. Times interest earned. b. Compare the liquidity, profitability, and solvency of the two companies.
Answer #1

Answer to Requirement a-1.

Target Corporation:
Current Ratio = Current Assets / Current Liabilities
Current Ratio = $11,573 / $12,777
Current Ratio = 0.91: 1

Wal-Mart Stores, Inc.:
Current Ratio = Current Assets / Current Liabilities
Current Ratio = $61,185 / $69,345
Current Ratio = 0.88: 1

Answer to Requirement a-2.

Target Corporation:
Accounts Receivable Turnover = Sales / Average Accounts Receivable
Accounts Receivable Turnover = $72,596 / $2,921
Accounts Receivable Turnover = 24.85 times

Wal-Mart Stores, Inc.:
Accounts Receivable Turnover = Sales / Average Accounts Receivable
Accounts Receivable Turnover = $476,294 / $6,723
Accounts Receivable Turnover = 70.85 times

Answer to Requirement a-3.

Target Corporation:
Average Collection Period = 365 / Accounts Receivable Turnover
Average Collection Period = 365 / 24.85
Average Collection Period = 14.69 days

Wal-Mart Stores, Inc.:
Average Collection Period = 365 / Accounts Receivable Turnover
Average Collection Period = 365 / 70.85
Average Collection Period = 5.15 days

Answer to Requirement a-4.

Target Corporation:
Inventory Turnover = Cost of Goods Sold / Average Inventory
Inventory Turnover = $51,160 / $8,335
Inventory Turnover = 6.14 times

Wal-Mart Stores, Inc.:
Inventory Turnover = Cost of Goods Sold / Average Inventory
Inventory Turnover = $358,069 / $44,331
Inventory Turnover = 8.08 times

Answer to Requirement a-5.

Target Corporation:
Days in Inventory = 365 / Inventory Turnover
Days in Inventory = 365 / 6.14
Days in Inventory = 59.45 days

Wal-Mart Stores, Inc.:
Days in Inventory = 365 / Inventory Turnover
Days in Inventory = 365 / 8.08
Days in Inventory = 45.17 days